A grocer has 2 kinds of nuts, one that costs $5 per kilogram and another that costs $5.40 per kilogram. How many kilograms of each type o fnut should be mixed in order to have 40 kg of a mixture worth $5.25 per kilogram?
Answer by checkley77(12569)

5.40x+5(40-x)=40*5.25
5.40x+200-5x=210
.4x=210-200
.4x=10
x=10/.4
x=25 kg of $5.40 nuts are needed.
40-25=15 kg of $ 5.00 nuts are needed.
Proof
5.40*25+5(40-25)=40*5.25
135+5*15=210
135+75=210
210=210